# Towards an Accessible, Noninvasive Micronutrient Status Assessment Method: A Comprehensive Review of Existing Techniques

### v1

@@ Appendix A Tables
Vitamin B9 conjuntiva pallor Central ridges redness in lips and tongue, swollen/inflammed/smooth tongue (glossitis), inflamed gums (gingivitis), dry/swollen/ulcerated lips (cheilosis), Aphthous Stomatitis (canker sores), inflamed/burning mouth, Atrophied papillae Anaemia, diabetes mellitus Correlated with fatigue
Vitamin B12 [5] conjuntiva pallor Pallor, clubbing (Koilonychia), transverse white lines (Muehrcke’s lines), excessive dryness, darkness in nails, curved nail ends, central ridges, longitudinal melanonychia Bilateral cracks/redness at corners of lips/mouth (angular cheilosis), swollen/inflammed/smooth tongue (glossitis), dry/swollen/ulcerated lips (cheilosis), pallor, Aphthous Stomatitis (canker sores), inflamed/burning mouth, Atrophied papillae, redness in lips and tongue, bleeding gums, tooth loss, tooth cavities Anaemia, Osteoporosis, sarcopenia Deficiency lowers HRV measurements, levels negitively correlated with sleep duration, levels positively correlated with sleep movement and self-assessed quality, night sweats, oxidative stress Correlated with fatigue Clinical symptoms can take years (typically 2-5) to appear because of storage levels, glossitis may present initally
Vitamin C Splinter hemmorage, excessive thinness, hapalonychia Intraoral mucosa and tongue inflammation, inflamed gums (gingivitis), thrush, tooth loss, tooth cavities Scurvy Supplementation improves blood pressure, helps regulate SNS Correlated with fatigue Deficiency can occur after 3-6 months of poor intake, signs of scurvy appear within 1 month of <10mg/day intake
Table 15: Physiological Symptoms of Micronutrient Deficiencies: Fat-Soluble Vitamins. Information from [123, 37, 39, 140, 142, 115, 17, 119]
@@
Vitamin D [144] Beau’s lines, longitudinal melanonychia, excessive thinness, hapalonychia inflamed gums (gingivitis) Cancer cachexia, cardiomyopathies/heart failure, Chronic obstructive pulmonary disease, osteoporosis, sarcopenia, critical to formation of hypocalcemia, depression Deficiency lowers HRV measurements, calcidiol deficiency lowers resting sympathovagal balance, calcitriol deficiency to worse reactions to stress, supplementation improves blood pressure
@@
Table 16: Physiological Symptoms of Micronutrient Deficiencies: Minerals. Information from [123, 37, 39, 140, 142, 115, 17, 119]
@@
Iron [58, 55] conjuntiva pallor, Redness/fissures in eyelid corners (Angular Palpebritis), conjunctiva redness/iritation, swollen/sticky eyelid, blue-tinted sclera Pallor, clubbing (Koilonychia), transverse white lines (Muehrcke’s lines), brittleness, excessive dryness, excessive thinness, darkness in nails, curved nail ends, central ridges, Onycholysis, onychorrhexis Bilateral cracks/redness at corners of lips/mouth (angular cheilosis), pallor, swollen/inflammed/smooth tongue (glossitis), Atrophied papillae, dry/swollen/ulcerated lips (cheilosis), thrush, inflamed/burning mouth, redness in lips and tongue Anaemia, cardiomyopathies/heart failure, osteoporosis Disrupts optimal function of endocrine and immune systems; positively correlated with sleep quality (disputed); IDA affects temp regulation and HRV (HRV disputed); low levels associated with higher HR Status has relation to energy levels and fatigue according to some sources; critical to oxygen binding; weakness; impaired cognitive function Multiple phases: depletion of stores (mild deficiency, can take several months), iron-deficiency erythropoiesis (erythrocyte production), then iron deficiency anemia (IDA)
Copper conjuntiva pallor Anaemia, chronic obstructive pulmonary disease, fatty liver disease, osteoporosis negitively correlated with sleep quality, reported to help regulate sns Abnormal lipid metabolism Some weeks to develop and not readily recognized, Usually manifests in acute conditions
Zinc [130] Conjunctiva inflamation/grittines (angular blepharitis) Beau’s lines, onychorrhexis, leukonychia, brittleness Changes in taste (inconsistently observed), dryness (Xerostomia), inflamed gums (gingivitis) Alcoholic hepatitis, cancer cachexia, chronic obstructive pulmonary disease, obesity, osteoporosis, sarcopenia, increased pneumonia risk Deficiency linked to increased blood pressure, positively correlated with sleep quality, reported to help regulate SNS, critical to ANS functionality according to some sources Light evidence of relationship between low dietary zinc and un-ideal metabolic response, correlated with fatigue Symptoms after "several months of low levels"
Iodine clubbing (Koilonychia) Goiter, hypothyroidism Critical to metabolic function Hypothyroidism occurs when intake falls below 10-20 mcg/d, goiter appears fairly quickly
Selenium excessive dryness, excessive thinness, pallor Cardiomyopathies/ heart failure, chronic obstructive pulmonary disease, obesity Intake reduces hypertrophy and oxidative stress, negitively effects blood pressure Effects on metabolism
@@
Table 17: Gold-Standard Methods of Assessing Micronutrient Imbalance: Water-Soluble Vitamins. Information from [69, 189, 115, 17, 119]
Detection Standard
Micronutrient Method Biomarker Matrix Intervals Impact of Inflammation
Vitamin B1 Erythrocyte transketolase activity coefficient assay Increase in erythrocyte transketolase activity Washed red blood cells Deficient: >25%; Insufficient: 15-25%; Sufficient: <15% None on direct plasma levels
Vitamin B2 Erythrocyte glutathione reductase activity coefficient assay Increase in erythrocyte glutathione reductase activity Washed red blood cells Deficient: >40%; Insufficient: 20-40%; Sufficient: <20% Decrease in plasma levels (erythrocyte assays are more stable)
Vitamin B3 LC-MS/MS Niacin metabolites (NMN and 2-pyr, limited representation of stores and recent intake) Urine Deficient: <5.8; Insufficient: 17.5-5.8; Sufficient: >17.5 mcmol/day Lack of evidence
Vitamin B5 LC-MS/MS Pantothenic acid (requires enzyme pretreatment) Whole blood Deficient: <0.22; Sufficient: 0.35-0.59 mg/L Lack of evidence
Vitamin B6 HPLC or LC-MS/MS pyridoxal phosphate (PLP) Plasma or serum Sufficient: >4.94 or >7.41 mcg/L plasma PLP (varies by source) Decrease in plasma PLP, no effect on RBC concentration
Vitamin B7 [44] LC-MS/MS for biotin; gel densitometry for MCC/PCC [107] Biotin (less sensitive); holo-MCC and holo-PCC (only reliable markers) Urine for biotin; WBCs for MCC/PCC Sufficient: 4.4-31 mcg/day urinary biotin, 8.2 arbitrary units holo-MCC, 9.1 arbitrary units holo-PCC None on biomarkers
Vitamin B9 LC-MS/MS Folate Serum for altered exposure and recent intake; Red blood cell for long term/3 month status and storage levels Sufficient: >3 ng/mL serum, >140 ng/mL RBC Lack of evidence
Vitamin B12 [5] GC-MS B12, confirmed with methylmalonic acid (MMA, also related to B2, B6, folate); no single ’gold standard’ Plasma or serum Deficient: <200-250 pg/mL B12, >0.03 mg/L MMA (some debate over this) Association with increased B12 levels
Vitamin C HPLC Ascorbate Plasma (some claim serum should be avoided) Deficient: 1.94; Insufficient: 2.11-4.05; Sufficient: 4.05 mg/L Decrease in plasma ascorbic acid (rapid, decrease when CRP >10 mg/L, normal values not detected if CRP >40 mg/L)
Table 18: Gold-Standard Methods of Assessing Micronutrient Imbalance: Fat-Soluble Vitamins. Information from [69, 189, 115, 17, 119]
Detection Standard
Micronutrient Method Biomarker Matrix Intervals Impact of Inflammation
Vitamin A LC-MS/MS Retinol (only sensitive to deficiency or excess in storage, affected by infection and protein/zinc deficiency). Best method is to indirectly measure reserves in liver over several days of administration Plasma or serum Severely deficient: <0.1; Deficient: 0.1-0.2; Sufficient: 0.3-1; Toxic: >1 mg/L retinol Decrease in serum retinol (adjustment equations exist but are not universally applicable, BRINDA R package)
Vitamin D [144] LC-MS/MS 25(OH)D (calcidiol) Plasma or serum Deficient: <12; Insufficient: 12-<20; Sufficient: 20-50; Toxic: >50 ng/mL (not definitively established/linked to clinical outcomes, varies based on assay and lab) Decrease in plasma levels (all values below reference ranges with CBP >40 mg/L)
Vitamin E LC-UV Ratio of Vit E to total blood lipids Plasma or serum Insufficient: <5.17 mg/L Vit E, <0.8 mg Vit E/g total lipid; Sufficient: 8.6-13 mg/L Vit E (adults have higher levels) Some effects (blood concentrations less interpretable at CRP >80 mg/L)
Vitamin K Immuno-based assays Plasma phylloquinone (usually for short term intake, no single ’gold standard’); prothrombin time (time to blood clot, only clinically relevant measure); variety of other ’functional’ biomarkers Plasma Deficient: <0.15; Sufficient (fasting): 0.15-1 mcg/L Status associated with lower inflammatory marker concentration
Table 19: Gold-Standard Methods of Assessing Micronutrient Imbalance: Minerals. Information from [69, 189, 115, 17, 119]
Detection Standard
Micronutrient Method Biomarker Matrix Intervals Impact of Inflammation
Iron [58, 55] Electrochemiluminescence immunoassay (ECLIA) Ferritin for deficiency (first phase, evaluates storage, inflated by infection); Iron increase after supplementation for malabsorption; Hemoglobin used to confirm IDA Serum Anemia: <10, Deficiency: 10-30 mcg/L ferritin Ferritin may be inflated, falsely normal/misleading (adjustment equations exist but are not universally applicable, BRINDA R package)
Copper ICP-MS Copper or ceruloplasmin (CP), neither reliable Serum Depletion: <50.8 (copper); Deficient: 50.8-76.2 (copper, high CRP); Sufficient: 63.5-158.9 mcg/dL (copper), 180-400 mg/L (CP) Increase in plasma concentrations
Zinc [130] Atomic Absorption Spectroscopy (AAS) Zinc (cut in half by SIRS, can be normal with clinical symptoms present, levels vary with time of day so it is recommended that albumin and CRP changes are taken into account) Plasma or serum Deficient: 70 women, 74 men; Insufficient: 70/74-80; Sufficient: 80-120 mcg/dL Decrease in plasma levels (significant when CRP exceeds 20 mg/L, adjustment equations exist but are not universally applicable, BRINDA R package)
Iodine ICP-MS Iodine Urine (24h or random), serum less recommended Depletion: <20, NA; Deficient: 20-100, <40; Sufficient: 100-300 mcg/24hr urine, 40-100 mcg/L serum (levels should be higher in those who are pregnant or lactating) Lack of evidence
Selenium AAS, ICP-MS Selenium (recent intake) or Selenoprotein P Plasma or serum Deficient: <60; Sufficient: >60; Toxicity: >474 to 948 mcg/L Se (intervals vary by source and population: women and black people have naturally lower concentrations) Decrease in plasma levels proportional to inflammation, can be adjusted for
Magnesium AAS Magnesium Serum (little correlation with overall status or tissue stores) and urine (after supplementation) Deficient: <18.23; Sufficient: 18.23-23.1 mg/L serum Mg

### v2

@@ Appendix A Tables
Vitamin B9 conjuntiva pallor Central ridges redness in lips and tongue, swollen/inflamed/smooth tongue (glossitis), inflamed gums (gingivitis), dry/swollen/ulcerated lips (cheilosis), Aphthous Stomatitis (canker sores), inflamed/burning mouth, Atrophied papillae Anemia, diabetes mellitus Correlated with fatigue
Vitamin B12 (Ada’s Medical Knowledge Team, 2023) conjuntiva pallor Pallor, clubbing (Koilonychia), transverse white lines (Muehrcke’s lines), excessive dryness, darkness in nails, curved nail ends, central ridges, longitudinal melanonychia Bilateral cracks/redness at corners of lips/mouth (angular cheilosis), swollen/inflamed/smooth tongue (glossitis), dry/swollen/ulcerated lips (cheilosis), pallor, Aphthous Stomatitis (canker sores), inflamed/burning mouth, Atrophied papillae, redness in lips and tongue, bleeding gums, tooth loss, tooth cavities Anemia, Osteoporosis, sarcopenia Deficiency lowers HRV measurements, levels negatively correlated with sleep duration, levels positively correlated with sleep movement and self-assessed quality, night sweats, oxidative stress Correlated with fatigue Clinical symptoms can take years (typically 2-5) to appear because of storage levels, glossitis may present initially
Vitamin C Splinter hemorrhage, excessive thinness, hapalonychia Intraoral mucosa and tongue inflammation, inflamed gums (gingivitis), thrush, tooth loss, tooth cavities Scurvy Supplementation improves blood pressure, helps regulate SNS Correlated with fatigue Deficiency can occur after 3-6 months of poor intake, signs of scurvy appear within 1 month of ¡10mg/day intake
Table 17. Physiological Symptoms of Micronutrient Deficiencies: Fat-Soluble Vitamins. Information from (of Nutrition and Dietetics, 2017; De and Seshadri, 2012; DiBaise and Tarleton, 2019; Radler and Lister, 2013; Reber et al., 2019; Mueller, 2017; Berger et al., 2022; National Institutes of Health, datea)
@@
Vitamin D (Ross and Institute of Medicine (2011), U. S.) Beau’s lines, longitudinal melanonychia, excessive thinness, hapalonychia inflamed gums (gingivitis) Cancer cachexia, cardiomyopathies/heart failure, Chronic obstructive pulmonary disease, osteoporosis, sarcopenia, critical to formation of hypocalcemia, depression Deficiency lowers HRV measurements, calcidiol deficiency lowers resting sympathovagal balance, calcitriol deficiency to worse reactions to stress, supplementation improves blood pressure
@@
Table 18. Physiological Symptoms of Micronutrient Deficiencies: Minerals. Information from (of Nutrition and Dietetics, 2017; De and Seshadri, 2012; DiBaise and Tarleton, 2019; Radler and Lister, 2013; Reber et al., 2019; Mueller, 2017; Berger et al., 2022; National Institutes of Health, datea)
@@
Iron (Gloria, date; Gerber, 2024) conjuntiva pallor, Redness/fissures in eyelid corners (Angular Palpebritis), conjunctiva redness/irritation, swollen/sticky eyelid, blue-tinted sclera Pallor, clubbing (Koilonychia), transverse white lines (Muehrcke’s lines), brittleness, excessive dryness, excessive thinness, darkness in nails, curved nail ends, central ridges, Onycholysis, onychorrhexis Bilateral cracks/redness at corners of lips/mouth (angular cheilosis), pallor, swollen/inflamed/smooth tongue (glossitis), Atrophied papillae, dry/swollen/ulcerated lips (cheilosis), thrush, inflamed/burning mouth, redness in lips and tongue Anemia, cardiomyopathies/heart failure, osteoporosis Disrupts optimal function of endocrine and immune systems; positively correlated with sleep quality (disputed); IDA affects temp regulation and HRV (HRV disputed); low levels associated with higher HR Status has relation to energy levels and fatigue according to some sources; critical to oxygen binding; weakness; impaired cognitive function Multiple phases: depletion of stores (mild deficiency, can take several months), iron-deficiency erythropoiesis (erythrocyte production), then iron deficiency anemia (IDA)
Copper Conjuntiva pallor Anemia, chronic obstructive pulmonary disease, fatty liver disease, osteoporosis negatively correlated with sleep quality, reported to help regulate sns Abnormal lipid metabolism Some weeks to develop and not readily recognized, Usually manifests in acute conditions
Zinc (Pat Bass, 2023) Conjunctiva inflammation/grittinesss (angular blepharitis) Beau’s lines, onychorrhexis, leukonychia, brittleness Changes in taste (inconsistently observed), dryness (Xerostomia), inflamed gums (gingivitis) Alcoholic hepatitis, cancer cachexia, chronic obstructive pulmonary disease, obesity, osteoporosis, sarcopenia, increased pneumonia risk Deficiency linked to increased blood pressure, positively correlated with sleep quality, reported to help regulate SNS, critical to ANS functionality according to some sources Light evidence of relationship between low dietary zinc and un-ideal metabolic response, correlated with fatigue Symptoms after ”several months of low levels”
Iodine clubbing (Koilonychia) Goiter, hypothyroidism Critical to metabolic function Hypothyroidism occurs when intake falls below 10-20 $\mu$g/d, goiter appears fairly quickly
Selenium excessive dryness, excessive thinness, pallor Cardiomyopathies/ heart failure, chronic obstructive pulmonary disease, obesity Intake reduces hypertrophy and oxidative stress, negatively effects blood pressure Effects on metabolism